OpenStudy (sphott51):

solve (x) / (x-4) - (1) / (x+3) = (28) / (x^2 -x -12) the answer is {-6} needs to show the work and idk how to get there

OpenStudy (anonymous):

\[\frac{x }{ (x-4) } - \frac{ 1 }{ (x+3) } = \frac{ 28 }{ (x^2-x-12) }\] is this your problem

OpenStudy (anonymous):

For the left side find a common denominator, remember when doing fractions \[\frac{ a }{ b } \pm \frac{ c }{ d } = \frac{ ad \pm bc }{ bd }\]

OpenStudy (anonymous):

What do you get when you factor \[(x^2-x-12)\]

OpenStudy (anonymous):

Hey are you there?

OpenStudy (sphott51):

Yeah, sorry, Wasn't on

OpenStudy (sphott51):

I got (x+4) (x-3)

OpenStudy (mathstudent55):

Close, but it's \((x - 4)(x + 3) \)

OpenStudy (sphott51):

Okay, I got it corrected now

OpenStudy (mathstudent55):

That means you have now: \(\dfrac{x}{x - 4} - \dfrac{1}{x + 3} = \dfrac{28}{(x - 4)(x + 3)}\) Ok?

OpenStudy (mathstudent55):

Now you need to subtract the fractions on the left side. Use the common denominator (x - 4)(x + 3) to do the subtraction.

OpenStudy (sphott51):

I subtract both numerator and denominator with (x-4)(x+3) ?

OpenStudy (mathstudent55):

When you subtract fractions with a common denominator, you use the same denominator and subtract the numerators.

OpenStudy (mathstudent55):

First, you need to multiply each fraction to get the common denominator.

OpenStudy (sphott51):

Like this?